A 83.1 kg skier rides a 1376.8 m long lift to the top of a mountain. The lift makes an angle of 23.9o with the horizontal. How much potential energy does she have at the top of the mountain?
1
Expert's answer
2020-02-10T09:21:56-0500

The  potential energy:


"E=mgh=mgl\sin{23.9\degree}"

"E=(83.1)(9.8)(1376.8)\sin{23.9\degree}=454260\ J"